OpenStudy (anonymous):

I need help on task 2 pg 67 algebra 1

OpenStudy (anonymous):

heres the problem:

OpenStudy (anonymous):

@Mokeira

OpenStudy (anonymous):

The top table, based on what it has given you, you need to find the equation. It is linear, so you need to use y=mx+b. That equation, should help you find when x+5, 6, and 7. Try that, if you want more explaination, I will give it. :) It's important to try to work these out yourself as much as you can.

OpenStudy (anonymous):

its the side table that they are talking about @dezmwas24

OpenStudy (mokeira):

yeah...finding the equation will make it easier

OpenStudy (anonymous):

whats a gp @Mokeira

OpenStudy (mokeira):

i mean AP...sorry!!!

OpenStudy (mokeira):

GP- geometric progression AP-Arithmetic progression

OpenStudy (anonymous):

I could be wrong, but it might be going up exponentially. What have you covered in the chapter?

OpenStudy (anonymous):

I thought it was going up by x2

OpenStudy (mokeira):

i thought we were adding multiples of 2

OpenStudy (anonymous):

yah that's what I meant

OpenStudy (mokeira):

so what would y be when x=5

OpenStudy (anonymous):

am I on the right track

OpenStudy (mokeira):

i think you are..i would do the exact same thing

OpenStudy (anonymous):

ok so now I simplify the numbers by adding

OpenStudy (anonymous):

Your reasoning is right, so yeah, you're on the right track. You still have to right an equation, right?

OpenStudy (mokeira):

yes...that looks like part (b)

OpenStudy (mokeira):

we still need an equation :( lol....

OpenStudy (anonymous):

yes @dezmwas24 but first I have to add the y values

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

Right.

OpenStudy (mokeira):

hmm....the equation is also super easy to get...shouldnt be hard. I was shying away from trying

OpenStudy (anonymous):

so now I have this

OpenStudy (mokeira):

have you added already?

OpenStudy (anonymous):

yes is that right

OpenStudy (anonymous):

Do you know how to find the equation?

OpenStudy (anonymous):

no

OpenStudy (mokeira):

i think you have sent the wrong one...but i am sure you are right...so now let us get the equation

OpenStudy (anonymous):

sorry

OpenStudy (mokeira):

you are right...how do we get the equation @dezmwas24

OpenStudy (anonymous):

??

OpenStudy (mokeira):

@dezmwas24

OpenStudy (anonymous):

Sorry. @247 Have you written an exponential equation before?

OpenStudy (anonymous):

I don't think so what is that

OpenStudy (anonymous):

ohh wait yes I have

OpenStudy (anonymous):

Does this formula look familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Your y values, if you noticed, they don't go up by the same amount every time. It goes up by 2, then 4, then 8, and so on. I think that's what makes it an exponential function.

OpenStudy (anonymous):

ok but I have no idea how to find the equation

OpenStudy (anonymous):

@JoannaBlackwelder @Mashy @ganeshie8

OpenStudy (anonymous):

@Mokeira

OpenStudy (anonymous):

I am so sorry. It has been a bit since I've done this, but I will try to figure it out.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@Gman344 @J.F @kirbykirby @inkyvoyd

OpenStudy (kirbykirby):

\[\Large y = \sum_{x=1}^n2^{x-1} \]

OpenStudy (kirbykirby):

You can use the formula for the sum of a geometric progression to find the value for x=15, or just sum it by hand

OpenStudy (anonymous):

@kirbykirby That's for the sum, not for an exponential equation.

OpenStudy (kirbykirby):

but \(y\) is finding a sum of terms when x = 3, y is \(2^0+2^1+2^2\)

OpenStudy (kirbykirby):

y is doing a sum of exponential functions if you want to view it that way

OpenStudy (anonymous):

So that equation you wrote earlier was the answer? I think I get it now.

OpenStudy (anonymous):

@247 Did you see what kirbykirby put?

OpenStudy (kirbykirby):

yes you can use that, but also a more efficient formula is using the formula for the sum of a geometric progression: \[\large \sum_{x=1}^nar^{x-1}=a\left( \frac{1-r^n}{1-r}\right) \] In this specific example we have \[\large \sum_{x=1}^n(1)(2)^{x-1}= \frac{1-2^n}{1-2} \]

OpenStudy (phi):

For part (b), your list looks good. There is an interesting pattern you may not notice in your numbers x y y+1 1 1 2 2 3 4 3 7 8 4 15 16 5 31 32 6 63 64 7 127 128 do you see that y+1 is a power of two? we can use that to find a formula to predict y when we know x

OpenStudy (phi):

to make it more clear x y y+1 y+1 1 1 2 2 2 3 4 2*2 3 7 8 2*2*2 4 15 16 2*2*2*2 5 31 32 2*2*2*2*2 6 63 64 2*2*2*2*2*2 7 127 128 2*2*2*2*2*2*2

OpenStudy (phi):

the idea is find y+1 from x (knowing it is a power of 2), then subtract 1 to get y

OpenStudy (phi):

That will be your answer for part C

OpenStudy (anonymous):

so @phi what will y be when x=15

OpenStudy (phi):

to find y when x is 15, you first need the formula Did you figure out the formula?

OpenStudy (anonymous):

no I don't know it @phi

OpenStudy (phi):

This is from yesterday x y y+1 y+1 1 1 2 2 2 3 4 2*2 3 7 8 2*2*2 4 15 16 2*2*2*2 5 31 32 2*2*2*2*2 6 63 64 2*2*2*2*2*2 7 127 128 2*2*2*2*2*2*2 any idea what the rule is for finding y+1 when x is 8 ?

OpenStudy (anonymous):

no

OpenStudy (phi):

Look at each x value, and then at how to compute y+1 for that x value what is the pattern ?

OpenStudy (anonymous):

I don't understand @phi

OpenStudy (phi):

are you saying you can't find a pattern in this: x y+1 1 2 2 2*2 3 2*2*2 4 2*2*2*2 5 2*2*2*2*2 6 2*2*2*2*2*2 7 2*2*2*2*2*2*2

OpenStudy (anonymous):

I don't have those numbers I have these @phi

OpenStudy (anonymous):

@Mokeira

OpenStudy (phi):

Yes, I know what numbers you have. Put that aside for the moment. Can you find the pattern in x y+1 1 2 2 2*2 3 2*2*2 4 2*2*2*2 5 2*2*2*2*2 6 2*2*2*2*2*2 7 2*2*2*2*2*2*2

OpenStudy (anonymous):

yes you add a *2 everytime

OpenStudy (phi):

be more specific. for example, when x=3, what is the rule for (y+1) ?

OpenStudy (phi):

This is close to being answered, if you try a little...

OpenStudy (anonymous):

\[2^{3}\] @phi

OpenStudy (phi):

yes, and if we generalize to x=n, y+1 = 2^n now back to your problem x y y+1 y+1 y+1 1 1 2 2 2^1 2 3 4 2*2 2^2 3 7 8 2*2*2 2^3 4 15 16 2*2*2*2 2^4 5 31 32 2*2*2*2*2 2^5 6 63 64 2*2*2*2*2*2 2^6 7 127 128 2*2*2*2*2*2*2 2^7 your list of numbers is the list labeled 'y' , the second column how do we change y+1 to get y ?

OpenStudy (anonymous):

I don't get where your getting these numbers @phi

OpenStudy (anonymous):

*2

OpenStudy (phi):

I noticed if we add 1 to your numbers (labeled "y" in the table) you get the numbers in the y+1 column, which (if you know your powers of 2) are powers of 2 It's easy to find the rule to find the numbers in the y+1 column if x is n , the number in the y+1 column is 2^n to get your number, subtract 1 so use 2^n - 1 as your rule.

OpenStudy (anonymous):

ok